Sally invested some money at 5% interest. sally also invested $66 more than three times that amount at 8%. how much is invested at each rate if so he receives $547.29 in interest after one year.
Answer by stanbon(75887)

Sally invested some money at 5% interest. sally also invested $66 more than three times that amount at 8%. how much is invested at each rate if so he receives $547.29 in interest after one year.
-----------------------
Interest + Interest = 547.29
0.05x + 0.08(3x+66) = 547.29
-----
5x + 24x + 8*66 = 54729
-------------------------------
29x = 54201
----
x = $1869 (amt invested at 5%)
3x+66 = $5673 (amt invested at 8%)
----------------
Cheers,
Stan H.
------------------